OpenStudy (anonymous):

what is the difference between writing 12 sin (pi/3 x+3)+12 and 12 sin(pi3 x +3) + 12 ?

OpenStudy (anonymous):

meant pi/3

OpenStudy (freckles):

you mean pi3 is pi/3?

OpenStudy (freckles):

if that is what you meant then I don't how what you wrote is any different and I'm not talking about just in meaning For example: If someone asked me if the following two sentences were the same: A dog ate a cat. A dog ate a cat. I would say these were the same sentence.

OpenStudy (freckles):

Because it is the same thing word by word.

OpenStudy (freckles):

\[12 \sin(\frac{\pi}{3}x+3)+12=12 \sin(\frac{\pi}{3}x+3)+12 \text{ for all } x\]